# Dynamic Gap: Safe Gap-based Navigation in Dynamic Environments

### v2

@@ I Introduction
Collision avoidance is of utmost importance for safe robot navigation. This task is typically handled by a local planner which utilizes sensory information to evade obstacles. One family of local planners is the gap-based planner [1], which identifies passable regions, or “gaps”, and synthesizes motion commands through them. With this emphasis on free space, gap-based planners are an approach based on the affordances of the environment [2], and they have shown great promise with capabilities of respecting dynamic, visual, and geometric constraints [3, 4, 5, 6, 7] as well as generating provably collision-free trajectories [8].
Despite this success, gap-based planners have only been extended to handling dynamic obstacle avoidance very recently [9], a challenge that accurately reflects the unknown, changing environments of the real world. Reactive planners designed for static environments are often deployed in dynamic environments, relying on sufficient runtimes to adapt to the evolving environment. However, explicitly accounting for dynamic obstacles during planning allows for more assured collision avoidance behaviors.
@@
This paper details an extension to prior work on the potential gap planner [8] involving augmentations to the planning framework to extend formal guarantees for safe navigation to dynamic environments under particular assumptions. Additional modules are also included to enable a more robust translation of planning success to non-ideal environments. This extension is referred to hereafter as dynamic gap and can be visualized in Figure 1. Egocentric free space tracking is integrated to predict how gaps will evolve over time, and guidance law-based policies are employed to synthesize collision-free trajectories that pass through moving gaps under ideal conditions. Additional modules are then applied to robustify planning performance in non-ideal conditions.
@@
This planner is open-sourced [^1] https://github.com/ivaROS/DynamicGap within the Arena-Rosnav benchmarking environment [10, 11, 12, 13] and available to test.
@@ II-A Perception Space and Gap-based Navigation
Most motion planners [14, 15, 16] opt to plan using Cartesian world frame environment representations such as cost maps or voxel grids. These approaches contrast with the perception-space approach to planning which involves keeping sensory input in its raw egocentric form to take advantage of the computational benefits that come with foregoing intensive data processing. All local planning steps downstream are then cast as ego-centric decision making processes.
Gap-based planners [3, 4, 5, 6, 17, 7, 18, 1, 19, 20, 9, 21, 22] are an example of perception space-based planners. They detect regions of collision-free space defined by either leading or trailing edges of obstacles which can be viewed as an alternative way of discretizing the environment, here in the egocentric polar space as opposed to common methods such as occupancy maps. These planners then synthesize local trajectories or reactive control inputs through these gaps. Some attention has been given to gaps in dynamic environments [22, 9], but these methods do not develop their theory through a perception-informed approach, instead opting to use ground truth agent pose information. In the proposed work, explicit attention is paid towards how the dynamics of gaps must be ascertained from scan data in order to understand how the local gaps evolve over time.
@@ II-B Guidance Laws
Guidance laws comprise a set of kinematic equations and feedback control laws that define collision course behavior between a pursuer and a target. While commonly affiliated with older forms of missile guidance, these laws have also seen use in many robotics applications [23, 24, 25].
Among the more established guidance laws, the two geometrical rules of pure pursuit and parallel navigation are the most popular. The pure pursuit rule, sometimes referred to as pursuit guidance, has the pursuer direct their velocity vector towards the target at all times, always keeping the target within the pursuer’s line of sight. Pure pursuit has seen a great deal of attention due to its simplicity [26, 27, 28], but this guidance law only leads to a collision if the pursuer is capable of traveling at a speed faster than that of the target.
The parallel navigation, or constant bearing, rule [29, 30] has the pursuer direct their velocity vector such that the direction of the line of sight between the pursuer and target remains constant while the distance between them decreases. This geometrical rule is capable of yielding collision course conditions even if the pursuer is traveling slower than the target. Furthermore, for a non-maneuvering target, meaning a target that is not changing its speed nor its heading direction, parallel navigation is the optimal guidance law which yields a minimum intercept time.
@@ III-A1 Gap Detection and Simplification
As input to the planner, we assume access to a 360°$\degree$ laser scan $\mathcal{L}$. Gap detection involves the parsing of this scan $\mathcal{L}$ to obtain a set of detected gaps $\mathcal{G}_{\rm det}$ that describe the instantaneous free space of the local environment, further details regarding how this gap detection policy is defined can be found at [8]. Once this set of gaps has been extracted from the laser scan, an additional pass through $\mathcal{G}_{\rm det}$ is performed to remove any redundant gaps and potentially merge adjacent gaps together, yielding a set of simplified gaps $\mathcal{G}_{\rm simp}$.
@@ III-A2 Gap Association
The association step is represented as a rectangular assignment problem, where the cost is equivalent to the distance between points across consecutive steps, $P_{\rm simp}^{t-1}$ and $P_{\rm simp}^{t}$. This assignment problem is then solved with the Hungarian Algorithm [31], producing a minimum distance mapping between $P_{\rm simp}^{t-1}$ and $P_{\rm simp}^{t}$. If the distance between two associated points exceeds a threshold $\tau_{assoc}$, then that point-to-point association is discarded.
@@ III-B Gap Propagation
    The gap closes, meaning if a gap that is shrinking over time reaches an angular span of 00 rad, and
@@ III-B1 Crossing condition
Figure 3: (a) Crossing condition for a gap, where between timesteps $t-1$ and $t$, the gap’s angular span reaches 00 rad. (b) Overlapping condition for a gap, where the gap’s angular span reaches $2\pi$ rad.
@@ III-C1 Guidance Law Analysis
We employ the parallel navigation policy [32], for which we assume a constant gap goal point speed $v_{g}$ as well as a constant ego-robot speed $v_{e}$. This policy then dictates the bearing $\theta_{e}$ in which $v_{e}$ will be applied. With relation to Figure 4, the parallel navigation policy is defined as $\dot{\beta}_{g}=0,\dot{r}_{g}<0$, where
@@ III-C3 Proof of Collision-Free Passage
[…]
@@ III-D Trajectory Scoring
The trajectory cost formulation is adapted from [8], with one key different: pose-wise scoring requires a laser scan for each timestep $t$ along the trajectory. We do not have direct access to this information, so we propagate gaps forward in time and back out propagated scans in practice.
@@ III-E2 Projection Operator
As a last resort safety filter to handle non-ideal circumstances including discrete time implementation and second-order dynamics, the proposed work also adapts the projection operator module from [8].
@@ IV-B Experiment 2: Assumption-violating Environments
To contextualize dynamic gap’s performance and understand the performance gap between ideal and non-ideal conditions, the planner is integrated into the Arena-Rosnav [10] benchmarking environment and compared against other state-of-the-art local planners. In this experiment, the classical motion planners DWA [33], TEB [15], and MPC [34] are tested along with the learned planners RLCA [35] and Trail [36]. Potential gap [8], the predecessor to dynamic gap, is also tested.
@@
Empty: Given that this environment is solely comprised of a single room, these trials can be used to evaluate each planner’s pure dynamic obstacle avoidance mechanisms. In this setting, the shortest path from start to goal is simply a straight line. Due to the lack of complexity in this setting, the variance in planning times is low, with most planners being able to reach the goal within $30-40$ seconds.
@@
[…]

### v3

@@ I Introduction
Collision avoidance is of utmost importance for safe robot navigation. This task is typically handled by a local planner which utilizes sensory information to evade obstacles. One family of local planners is the gap-based planner [mujahad_closest_2010], which identifies passable regions, or “gaps”, and synthesizes motion commands through them. With this emphasis on free space, gap-based planners are an approach based on the affordances of the environment [gibson_ecological_2014], and they have shown great promise with capabilities of respecting dynamic, visual, and geometric constraints [sezer_novel_2012, mujahed_admissible_2018, mujahed_safe_2013, mujahed_new_2016, mujahed_tangential_2013] as well as generating provably collision-free trajectories [xu_potential_2021].
Despite this success, gap-based planners have only been extended to handling dynamic obstacle avoidance very recently [contarli_predictive_2024], a challenge that accurately reflects the unknown, changing environments of the real world. Reactive planners designed for static environments are often deployed in dynamic environments, relying on sufficient runtimes to adapt to the evolving environment. However, explicitly accounting for dynamic obstacles during planning allows for more assured collision avoidance behaviors.
@@
This paper details an extension to prior work on the potential gap planner [xu_potential_2021] involving augmentations to the planning framework to extend formal guarantees for safe navigation to dynamic environments under particular assumptions. Additional modules are also included to enable a more robust translation of planning success to non-ideal environments. This extension is referred to hereafter as dynamic gap and can be visualized in Figure 1. Egocentric free space tracking is integrated to predict how gaps will evolve over time, and guidance law-based policies are employed to synthesize collision-free trajectories that pass through moving gaps under ideal conditions. Additional modules are then applied to robustify planning performance in non-ideal conditions.
@@
This planner is open-sourced [^1] https://github.com/ivaROS/DynamicGap within the Arena-Rosnav benchmarking environment [kastner_arena-rosnav_2021, kastner_arena-bench_2022, kastner_arena-rosnav_2023, kastner_arena_2024] and available to test.
@@ II-A Perception Space and Gap-based Navigation
Most motion planners [lavalle_rapidly-exploring_1998, rosmann_timed-elastic-bands_2015, connell_dynamic_2017] opt to plan using Cartesian world frame environment representations such as cost maps or voxel grids. These approaches contrast with the perception-space approach to planning which involves keeping sensory input in its raw egocentric form to take advantage of the computational benefits that come with foregoing intensive data processing. All local planning steps downstream are then cast as ego-centric decision making processes.
Gap-based planners [sezer_novel_2012, mujahed_admissible_2018, mujahed_safe_2013, mujahed_new_2016, feng_gpf-bg_2023, mujahed_tangential_2013, mujahed_admissible_2017, mujahad_closest_2010, ullah_fnug_2022, demir_improved_2017, contarli_predictive_2024, feng_safer_2023, chen_safe_2022] are an example of perception space-based planners. They detect regions of collision-free space defined by either leading or trailing edges of obstacles which can be viewed as an alternative way of discretizing the environment, here in the egocentric polar space as opposed to common methods such as occupancy maps. These planners then synthesize local trajectories or reactive control inputs through these gaps. Some attention has been given to gaps in dynamic environments [chen_safe_2022, contarli_predictive_2024], but these methods do not develop their theory through a perception-informed approach, instead opting to use ground truth agent pose information. In the proposed work, explicit attention is paid towards how the dynamics of gaps must be ascertained from scan data in order to understand how the local gaps evolve over time.
@@ II-B Guidance Laws
Guidance laws comprise a set of kinematic equations and feedback control laws that define collision course behavior between a pursuer and a target. While commonly affiliated with older forms of missile guidance, these laws have also seen use in many robotics applications [noauthor_implementation_nodate, wellhausen_artplanner_2023, fiorini_motion_1998].
Among the more established guidance laws, the two geometrical rules of pure pursuit and parallel navigation are the most popular. The pure pursuit rule, sometimes referred to as pursuit guidance, has the pursuer direct their velocity vector towards the target at all times, always keeping the target within the pursuer’s line of sight. Pure pursuit has seen a great deal of attention due to its simplicity [bernhart_polygons_1959, bernhart_curves_1959, bruckstein_why_1993], but this guidance law only leads to a collision if the pursuer is capable of traveling at a speed faster than that of the target.
The parallel navigation, or constant bearing, rule [rajasekhar_fuzzy_2000, ulybyshev_terminal_2005] has the pursuer direct their velocity vector such that the direction of the line of sight between the pursuer and target remains constant while the distance between them decreases. This geometrical rule is capable of yielding collision course conditions even if the pursuer is traveling slower than the target. Furthermore, for a non-maneuvering target, meaning a target that is not changing its speed nor its heading direction, parallel navigation is the optimal guidance law which yields a minimum intercept time.
@@ III-A1 Gap Detection and Simplification
[…]
@@ III-A2 Gap Association
The association step is represented as a rectangular assignment problem, where the cost is equivalent to the distance between points across consecutive steps, $P_{\rm simp}^{t-1}$ and $P_{\rm simp}^{t}$. This assignment problem is then solved with the Hungarian Algorithm [kuhn_hungarian_1955], producing a minimum distance mapping between $P_{\rm simp}^{t-1}$ and $P_{\rm simp}^{t}$. If the distance between two associated points exceeds a threshold $\tau_{assoc}$, then that point-to-point association is discarded.
@@ III-B Gap Propagation
    The gap closes, meaning if a gap that is shrinking over time reaches an angular span of 0 rad, and
@@ III-B1 Crossing condition
Figure 3: (a) Crossing condition for a gap, where between timesteps t1t-1 and $t$, the gap’s angular span reaches 0 rad. (b) Overlapping condition for a gap, where the gap’s angular span reaches $2\pi$ rad.
@@ III-C1 Guidance Law Analysis
We employ the parallel navigation policy [shneydor_missile_1998], for which we assume a constant gap goal point speed $v_{g}$ as well as a constant ego-robot speed $v_{e}$. This policy then dictates the bearing $\theta_{e}$ in which $v_{e}$ will be applied. With relation to Figure 4, the parallel navigation policy is defined as $\dot{\beta}_{g}=0,\dot{r}_{g}<0$, where
@@ III-C3 Proof of Collision-Free Passage
$\begin{split}\mathbf{p}_{g}&=\kappa\mathbf{p}_{l}+(1-\kappa)\mathbf{p}_{r},\\ \mathbf{v}_{g}&=\kappa\mathbf{v}_{l}+(1-\kappa)\mathbf{v}_{r},\hskip 14.22636pt\kappa\in[0,1].\end{split}$ (15)
@@ III-D Trajectory Scoring
The trajectory cost formulation is adapted from [xu_potential_2021], with one key different: pose-wise scoring requires a laser scan for each timestep $t$ along the trajectory. We do not have direct access to this information, so we propagate gaps forward in time and back out propagated scans in practice.
@@ III-E2 Projection Operator
As a last resort safety filter to handle non-ideal circumstances including discrete time implementation and second-order dynamics, the proposed work also adapts the projection operator module from [xu_potential_2021].
@@ IV-B Experiment 2: Assumption-violating Environments
To contextualize dynamic gap’s performance and understand the performance gap between ideal and non-ideal conditions, the planner is integrated into the Arena-Rosnav [kastner_arena-rosnav_2021] benchmarking environment and compared against other state-of-the-art local planners. In this experiment, the classical motion planners DWA [fox_dynamic_1997], TEB [rosmann_timed-elastic-bands_2015], and MPC [rosmann_online_2021] are tested along with the learned planners RLCA [long_towards_2018] and Trail [xie_drl-vo_2023]. Potential gap [xu_potential_2021], the predecessor to dynamic gap, is also tested.
@@
Empty: Given that this environment is solely comprised of a single room, these trials can be used to evaluate each planner’s pure dynamic obstacle avoidance mechanisms. In this setting, the shortest path from start to goal is simply a straight line. Due to the lack of complexity in this setting, the variance in planning times is low, with most planners being able to reach the goal within 304030-40 seconds.
@@
The ability of dynamic gap to predict the feasibility of local gaps before committing to them proved paramount in avoiding collisions in this environment. The proposed planner suffered no collisions over its trials, though this “patience” did result in longer planning times overall.